w + 3 = b
b x 2 = j
w + b + j = 41
I cant seem to find out how to get w,b and j. Any and all help is appreciated.

Answer by JoelSchwartz(130)

b=w+3
j=2b
2(w+3)=j
2w+6=j
w+b+j=41
w+w+3+2w+6=41
4w=41-9
4w=32
w=8
b=11
j=22